OpenStudy (gabylovesyou):

Write an equation of a line that passes through the point (2, 3) and is parallel to the line y = 3 over 2x + 5. y = negative 2 over 3x y = negative 2 over 3x − 7 y = 3 over 2x y = 3 over 2x + 7

OpenStudy (gabylovesyou):

@johnweldon1993

OpenStudy (johnweldon1993):

So first thing...parallel lines have ___ slopes?

OpenStudy (gabylovesyou):

same ?

OpenStudy (johnweldon1993):

Right...so the line we are looking for has the same slope as \[\large y = \frac{3}{2}x + 5\] so it would have the slope 3 over 2 right? now..we know a slope...and we know a point on the line, lets use the point slope form of a line to find out what the equation is \[\large y - y_0 = m(x - x_0)\] m is the slope.....x_o and y_o are the points we have...(2,3) so if we plug all that in...we have \[\large y - 3 = \frac{3}{2}(x - 2)\] can you simplify that for me :)

OpenStudy (gabylovesyou):

y = 3x/2

OpenStudy (johnweldon1993):

Perfect :)

OpenStudy (gabylovesyou):

its not in my choices ;/

OpenStudy (johnweldon1993):

It's the third answer choice :) \[\large y = \frac{3}{2}x\]

OpenStudy (gabylovesyou):

ohhhhhhhhhh gotcha ... thank you :3

OpenStudy (johnweldon1993):

Anytime :)